Find the distance of the point (2,5) from the line 3x+y+4=0 measured parallel to the line having slope 3/4

Grade:11

1 Answers

Arun
6 years ago
Dear Dharmik
 
The parallel line has slope m=(3/4) and passes by (2,5) 
y-5=(3/4) (x-2) 
4y-20= 3x-6 
4y =3x+14 
3x=4y-14 

This line intercepts to line 3x+y+4=0 when 
(4y-14) +y+4=0 
5y= 10 
y=2 
x=(8-14)/3 = -2 
thus, distance is from (2,5) to (-2,2) 

D= sqrt ( 2-(-2)) ^2 + (5-2)^2 

D= sqrt (16+9) = sqrt25 =5 
 
Regards
Arun (AskIITians forum expert)